36 grams of aspirin is burnt with a flame in presence of excess oxygen until nothing remains. The byproduct gases are collected and it is found that there are 79.2 g of pure CO2 as one of the products. How many carbon atoms are there in each formula unit of Aspirin (MW = 180g)?


1
Expert's answer
2021-10-20T02:52:10-0400

Mass of "CO_2=79.2g"

Therefore number of moles "=\frac{mass}{molar mass}= \frac{79.2g}{44g/mol}"


"=1.8mole CO_2"


1 mole "CO_2" contains only 1 mole C

Therefore 1.8 "CO_2" must contain 1.8 mole C

Mass of aspirin"=36g"

Molar mass"=180g\/mol"

Therefore moles of aspirin"=\frac{36}{180} =0.2moles"


Since total mole of C in 0.2 mole of aspirin = 1.8 mole this means in 1 mole aspirin there will be "\frac{1.8}{0.2}" "=9"

Hence aspirin must contains 9 C atoms